# Initial value space of the fourth order Painlevé system with $A_{5}^{(1)}\times A_{1}^{(1)}$ symmetry

### v2

@@ 3 Accessible singularities
In the previous section, we have shown that the system $\mathcal{H}$ has the singularities only on $W_{ij}\times B\ (i=0,1,2,\ j=1,2)$. On these $W_{ij}\times B$, there exist accessible singularities where solutions to the system $\mathcal{H}^{0}$ pass through. For example, the accessible singularities on $W_{01}\times B$ are given by the condition $p_{1}^{(01)}=F_{1}^{(01)}=F_{2}^{(01)}=G_{2}^{(01)}=0$. It follows from the form of system (2.6). In this section, we list all of the accessible singularities of the system $\mathcal{H}$.
@@
This lemma will be shown in Appendix A. In the following, we use a notation of the accessible singularities
@@
We have reduced the accessible singularities of the system $\mathcal{H}$ to
$\begin{split}&C_{1}^{(01)},\ C_{2}^{(01)},\ C_{3}^{(01)},\ C_{2}^{(02)},\ C_{3}^{(02)},\ C_{3}^{(11)},\ C_{1}^{(12)},\ C_{2}^{(12)},\ C_{3}^{(12)},\ C_{3}^{(21)},\ C_{1}^{(22)},\ C_{3}^{(22)},\\ &P_{3}^{(01)},\ P_{4}^{(01)},\ P_{6}^{(01)},\ P_{7}^{(01)},\ P_{3}^{(12)},\ P_{4}^{(12)},\ P_{6}^{(12)},\ P_{7}^{(12)},\ P_{3}^{(22)},\ P_{4}^{(22)},\ P_{6}^{(22)},\ P_{7}^{(22)}\end{split}$ (3.2)
thanks to Lemma 3.2 and Lemma 3.3. Let $C$ be a union of the accessible singularities defined by
$C=C_{1}^{(01)}\cup C_{2}^{(01)}\cup C_{3}^{(01)}\cup C_{2}^{(02)}\cup C_{3}^{(02)}\cup C_{3}^{(11)}\cup C_{1}^{(12)}\cup C_{2}^{(12)}\cup C_{3}^{(12)}\cup C_{3}^{(21)}\cup C_{1}^{(22)}\cup C_{3}^{(22)}.$
Then we obtain the following lemma.
@@
The union $C$ is described as a disjoint union of 6 connected components
$C=C_{2}^{(01)}\sqcup C_{2}^{(12)}\sqcup C_{2}^{(22)}\sqcup\left(C_{3}^{(01)}\cup C_{3}^{(02)}\right)\sqcup\left(C_{3}^{(11)}\cup C_{3}^{(12)}\right)\sqcup\left(C_{3}^{(21)}\cup C_{3}^{(22)}\right).$
Moreover, $C$ does not contain any accessible singularity $P_{j}^{(kl)}$ appeared in (3.2).
Proof.
We first consider the accessible singularities except for $C_{1}^{(01)},C_{1}^{(12)},C_{1}^{(22)}$. Using the homogeneous coordinates (2.2), we have
$C_{2}^{(01)}$ $=\left\{\left(\xi,\zeta^{(0)},t\right)\bigm{|}\xi_{0}-\xi_{1}=\zeta_{0}^{(0)}=\zeta_{2}^{(0)}=0,\ \xi_{0}\neq 0,\ \zeta_{1}^{(0)}\neq 0\right\}$
$=\left\{\left([\xi_{0}:\xi_{0}:\xi_{2}],[0:1:0],t\right)\bigm{|}\xi_{0}\neq 0\right\},$
$C_{2}^{(12)}$ $=\left\{\left(\xi,\zeta^{(1)},t\right)\bigm{|}\xi_{1}-\xi_{2}=\zeta_{0}^{(1)}=\zeta_{1}^{(1)}=0,\ \xi_{1}\neq 0,\ \zeta_{2}^{(1)}\neq 0\right\}$
$=\left\{\left([\xi_{0}:\xi_{1}:\xi_{1}],[0:0:1],t\right)\bigm{|}\xi_{1}\neq 0\right\},$
$C_{2}^{(22)}$ $=\left\{\left(\xi,\zeta^{(2)},t\right)\bigm{|}t\xi_{0}-\xi_{2}=\zeta_{0}^{(2)}=\zeta_{1}^{(2)}=0,\ \xi_{2}\neq 0,\ \zeta_{2}^{(2)}\neq 0\right\}$
$=\left\{\left(\left[t^{-1}\xi_{2}:\xi_{1}:\xi_{2}\right],[0:0:1],t\right)\bigm{|}\xi_{2}\neq 0\right\},$
$C_{3}^{(01)}$ $=\left\{\left(\xi,\zeta^{(0)},t\right)\bigm{|}\xi_{1}=\xi_{2}=\zeta_{0}^{(0)}=0,\ \xi_{0}\neq 0,\ \zeta_{1}^{(0)}\neq 0\right\}$
$=\left\{\left([1:0:0],\left[0:\zeta_{1}^{(0)}:\zeta_{2}^{(0)}\right],t\right)\bigm{|}\zeta_{1}^{(0)}\neq 0\right\},$
$C_{3}^{(02)}$ $=\left\{\left(\xi,\zeta^{(0)},t\right)\bigm{|}\xi_{1}=\xi_{2}=\zeta_{0}^{(0)}=0,\ \xi_{0}\neq 0,\ \zeta_{2}^{(0)}\neq 0\right\}$
$=\left\{\left([1:0:0],\left[0:\zeta_{1}^{(0)}:\zeta_{2}^{(0)}\right],t\right)\bigm{|}\zeta_{2}^{(0)}\neq 0\right\},$
$C_{3}^{(11)}$ $=\left\{\left(\xi,\zeta^{(1)},t\right)\bigm{|}\xi_{0}=\xi_{2}=\zeta_{0}^{(1)}=0,\ \xi_{1}\neq 0,\ \zeta_{1}^{(1)}\neq 0\right\}$
$=\left\{\left([0:1:0],\left[0:\zeta_{1}^{(1)}:\zeta_{2}^{(1)}\right],t\right)\bigm{|}\zeta_{1}^{(1)}\neq 0\right\},$
$C_{3}^{(12)}$ $=\left\{\left(\xi,\zeta^{(1)},t\right)\bigm{|}\xi_{0}=\xi_{2}=\zeta_{0}^{(1)}=0,\ \xi_{1}\neq 0,\ \zeta_{2}^{(1)}\neq 0\right\}$
$=\left\{\left([0:1:0],\left[0:\zeta_{1}^{(1)}:\zeta_{2}^{(1)}\right],t\right)\bigm{|}\zeta_{2}^{(1)}\neq 0\right\},$
$C_{3}^{(21)}$ $=\left\{\left(\xi,\zeta^{(2)},t\right)\bigm{|}\xi_{0}=\xi_{1}=\zeta_{0}^{(2)}=0,\ \xi_{2}\neq 0,\ \zeta_{1}^{(2)}\neq 0\right\}$
$=\left\{\left([0:0:1],\left[0:\zeta_{1}^{(2)}:\zeta_{2}^{(2)}\right],t\right)\bigm{|}\zeta_{1}^{(2)}\neq 0\right\},$
$C_{3}^{(22)}$ $=\left\{\left(\xi,\zeta^{(2)},t\right)\bigm{|}\xi_{0}=\xi_{1}=\zeta_{0}^{(2)}=0,\ \xi_{2}\neq 0,\ \zeta_{2}^{(2)}\neq 0\right\}$
$=\left\{\left([0:0:1],\left[0:\zeta_{1}^{(2)}:\zeta_{2}^{(2)}\right],t\right)\bigm{|}\zeta_{2}^{(2)}\neq 0\right\}.$
It implies
$\left(C_{3}^{(01)}\cup C_{3}^{(02)}\right)\cap\left(C_{3}^{(11)}\cup C_{3}^{(12)}\right)=\emptyset,$
$\left(C_{3}^{(01)}\cup C_{3}^{(02)}\right)\cap\left(C_{3}^{(21)}\cup C_{3}^{(22)}\right)=\emptyset,$
$\left(C_{3}^{(11)}\cup C_{3}^{(12)}\right)\cap\left(C_{3}^{(21)}\cup C_{3}^{(22)}\right)=\emptyset$

### v3

@@ 4 Accessible singularities
In the previous section, we have shown that the system $\mathcal{H}$ has the singularities only on $W_{ij}\times B\ (i=0,1,2,\ j=1,2)$. There exist accessible singularities on these $W_{ij}\times B$ where leaves corresponding to solutions to $\mathcal{H}$ passing through are transversal leaves. For example, the accessible singularities on $W_{01}\times B$ are given by the condition $p_{1}^{(01)}=F_{1}^{(01)}=F_{2}^{(01)}=G_{2}^{(01)}=0$. It follows from the form of system (3.6). In this section, we list all of the accessible singularities of $\mathcal{H}$.
@@
This lemma will be shown in Appendix A. In the following, we use a notation of the singularities
@@
We have reduced the accessible singularities of $\mathcal{H}$ to
$\begin{split}&C_{1}^{(01)},\ C_{2}^{(01)},\ C_{3}^{(01)},\ C_{3}^{(02)},\ C_{1}^{(11)},\ C_{3}^{(11)},\ C_{2}^{(12)},\ C_{3}^{(12)},\ C_{1}^{(21)},\ C_{3}^{(21)},\ C_{2}^{(22)},\ C_{3}^{(22)},\\ &P_{3}^{(01)},\ P_{4}^{(01)},\ P_{6}^{(01)},\ P_{7}^{(01)},\ P_{3}^{(12)},\ P_{4}^{(12)},\ P_{6}^{(12)},\ P_{7}^{(12)},\ P_{3}^{(22)},\ P_{4}^{(22)},\ P_{6}^{(22)},\ P_{7}^{(22)}\end{split}$ (4.2)
thanks to Lemma 4.2 and Lemma 4.3. Let $C^{0}$ and $P^{0}$ be unions of the accessible singularities defined by
$C^{0}=C_{1}^{(01)}\cup C_{2}^{(01)}\cup C_{3}^{(01)}\cup C_{3}^{(02)}\cup C_{1}^{(11)}\cup C_{3}^{(11)}\cup C_{2}^{(12)}\cup C_{3}^{(12)}\cup C_{1}^{(21)}\cup C_{3}^{(21)}\cup C_{2}^{(22)}\cup C_{3}^{(22)}$
and
$P^{0}=P_{3}^{(01)}\cup P_{4}^{(01)}\cup P_{6}^{(01)}\cup P_{7}^{(01)}\cup P_{3}^{(12)}\cup P_{4}^{(12)}\cup P_{6}^{(12)}\cup P_{7}^{(12)}\cup P_{3}^{(22)}\cup P_{4}^{(22)}\cup P_{6}^{(22)}\cup P_{7}^{(22)}$
respectively. Then we obtain the following lemma.
@@
The union $C^{0}$ is described as a disjoint union of three connected components
$C^{0}=C_{1}^{0}\sqcup C_{2}^{0}\sqcup C_{3}^{0},$
where
$C_{1}^{0}=C_{1}^{(01)}\cup C_{3}^{(01)}\cup C_{3}^{(02)}\cup C_{2}^{(12)},$
$C_{2}^{0}=C_{1}^{(11)}\cup C_{3}^{(11)}\cup C_{3}^{(12)}\cup C_{2}^{(22)},$
$C_{3}^{0}=C_{1}^{(21)}\cup C_{3}^{(21)}\cup C_{3}^{(22)}\cup C_{2}^{(01)}.$
Moreover, $C^{0}$ and $P^{0}$ do not intersect.
This lemma will be shown in Appendix B. Lemma 4.2, Lemma 4.3 and Lemma 4.4 imply the following proposition.
Proposition 4.5.
All of the accessible singularities of the system $\mathcal{H}$ is reduced to $C^{0}$ and $P^{0}$.
In the last, we investigate solutions passing through the singularity $D^{0}\setminus(C^{0}\cup P^{0})$. We consider a point $\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)},t\right)=(a_{0},b_{0},0,d_{0},t_{0})$ such that $F_{1}^{(01)}(a_{0},b_{0},0,d_{0},t_{0})\neq 0$ as an example. System (3.6) is rewritten to a Pfaff system
$F_{1}^{(01)}dt-t(t-1)p_{1}^{(01)}dq_{1}=0,$
$F_{1}^{(01)}dq_{2}-F_{2}^{(01)}dq_{1}=0,$
$F_{1}^{(01)}dp_{1}^{(01)}-p_{1}^{(01)}G_{1}^{(01)}dq_{1}=0,$
$F_{1}^{(01)}dp_{2}^{(01)}-G_{2}^{(01)}dq_{1}=0.$
Then the solution to this system passing through a point $\left(q_{1},q_{2},p_{1}^{(01)},p_{2}^{(01)},t\right)=(a_{0},b_{0},0,d_{0},t_{0})$ satisfies
$p_{1}^{(01)}=0,\quad t=t_{0}.$
This solution corresponds to a vertical leaf. We can show that all of the leaves corresponding to the solutions to $\mathcal{H}$ passing through the singularity $D^{0}\setminus(C^{0}\cup P^{0})$ are vertical leaves in a similar manner.
5 Resolving the accessible singularities $C^{0}$
In the previous section, we have determined the accessible singularities of the system $\mathcal{H}$. As will be seen below, there are infinitely many solutions passing through each accessible singularity. In this section, we resolve the accessible singularities $C^{0}$ by introducing new coordinates which parametrize these solutions. We give the result of this section in advance.
Proposition 5.1.
Coordinates resolving the accessible singularities $C^{0}$ are given by
$C_{1}^{(01)}:$ $\left(x_{1}^{[01]},y_{1}^{[01]},z_{1}^{[01]},w_{1}^{[01]}\right)=\left(-\{(q_{1}-q_{2})p_{1}-\alpha_{2}\}p_{1},q_{2},\dfrac{1}{p_{1}},p_{1}+p_{2}\right),$
$C_{3}^{(01)}:$ $\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)=\left(-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{1},q_{2}p_{1},\dfrac{1}{p_{1}},\dfrac{p_{2}}{p_{1}}\right),$
$\left(\widehat{x}_{3}^{[01]},\widehat{y}_{3}^{[01]},\widehat{z}_{3}^{[01]},\widehat{w}_{3}^{[01]}\right)=r_{2}\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right),$
$C_{3}^{(02)}:$ $\left(x_{3}^{[02]},y_{3}^{[02]},z_{3}^{[02]},w_{3}^{[02]}\right)=\left(q_{1}p_{2},-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{2},\dfrac{p_{1}}{p_{2}},\dfrac{1}{p_{2}}\right),$
$\left(\widehat{x}_{3}^{[02]},\widehat{y}_{3}^{[02]},\widehat{z}_{3}^{[02]},\widehat{w}_{3}^{[02]}\right)=r_{2}\left(x_{3}^{[02]},y_{3}^{[02]},z_{3}^{[02]},w_{3}^{[02]}\right),$
$C_{2}^{(12)}:$ $\left(x_{2}^{[12]},y_{2}^{[12]},z_{2}^{[12]},w_{2}^{[12]}\right)=\left(q_{1}^{(1)},-\left\{\left(q_{2}^{(1)}-1\right)p_{2}^{(1)}-\alpha_{2}\right\}p_{2}^{(1)},p_{1}^{(1)},\dfrac{1}{p_{2}^{(1)}}\right)$
for $C_{1}^{0}$,
$C_{1}^{(11)}:$ $\left(x_{1}^{[11]},y_{1}^{[11]},z_{1}^{[11]},w_{1}^{[11]}\right)=\left(-\left\{\left(q_{1}^{(1)}-\dfrac{q_{2}^{(1)}}{t}\right)p_{1}^{(1)}-\alpha_{4}\right\}p_{1}^{(1)},q_{2}^{(1)},\dfrac{1}{p_{1}^{(1)}},p_{2}^{(1)}+\dfrac{p_{1}^{(1)}}{t}\right),$
$C_{3}^{(11)}:$ $\left(x_{3}^{[11]},y_{3}^{[11]},z_{3}^{[11]},w_{3}^{[11]}\right)=\left(-\left(q_{1}^{(1)}p_{1}^{(1)}+q_{2}^{(1)}p_{2}^{(1)}-\alpha_{1}+\eta\right)p_{1}^{(1)},q_{2}^{(1)}p_{1}^{(1)},\dfrac{1}{p_{1}^{(1)}},\dfrac{p_{2}^{(1)}}{p_{1}^{(1)}}\right),$
$\left(\widehat{x}_{3}^{[11]},\widehat{y}_{3}^{[11]},\widehat{z}_{3}^{[11]},\widehat{w}_{3}^{[11]}\right)=r_{4}\left(x_{3}^{[11]},y_{3}^{[11]},z_{3}^{[11]},w_{3}^{[11]}\right),$
$C_{3}^{(12)}:$ $\left(x_{3}^{[12]},y_{3}^{[12]},z_{3}^{[12]},w_{3}^{[12]}\right)=\left(q_{1}^{(1)}p_{2}^{(1)},-\left(q_{1}^{(1)}p_{1}^{(1)}+q_{2}^{(1)}p_{2}^{(1)}-\alpha_{1}+\eta\right)p_{2}^{(1)},\dfrac{p_{1}^{(1)}}{p_{2}^{(1)}},\dfrac{1}{p_{2}^{(1)}}\right),$
$\left(\widehat{x}_{3}^{[12]},\widehat{y}_{3}^{[12]},\widehat{z}_{3}^{[12]},\widehat{w}_{3}^{[12]}\right)=r_{4}\left(x_{3}^{[12]},y_{3}^{[12]},z_{3}^{[12]},w_{3}^{[12]}\right),$
$C_{2}^{(22)}:$ $\left(x_{2}^{[22]},y_{2}^{[22]},z_{2}^{[22]},w_{2}^{[22]}\right)=\left(q_{1}^{(2)},-\left\{\left(q_{2}^{(2)}-\frac{1}{t}\right)p_{2}^{(2)}-\alpha_{4}\right\}p_{2}^{(2)},p_{1}^{(2)},\frac{1}{p_{2}^{(2)}}\right)$
for $C_{2}^{0}$ and
$C_{2}^{(01)}:$ $\left(x_{2}^{[01]},y_{2}^{[01]},z_{2}^{[01]},w_{2}^{[01]}\right)=\left(-\{(q_{1}-1)p_{1}-\alpha_{0}\}p_{1},q_{2},\dfrac{1}{p_{1}},p_{2}\right),$
$C_{1}^{(21)}:$ $\left(x_{1}^{[21]},y_{1}^{[21]},z_{1}^{[21]},w_{1}^{[21]}\right)=\left(-\left\{\left(q_{1}^{(2)}-q_{2}^{(2)}\right)p_{1}^{(2)}-\alpha_{0}\right\}p_{1}^{(2)},q_{2}^{(2)},\dfrac{1}{p_{1}^{(2)}},p_{1}^{(2)}+p_{2}^{(2)}\right),$
$C_{3}^{(21)}:$ $\left(x_{3}^{[21]},y_{3}^{[21]},z_{3}^{[21]},w_{3}^{[21]}\right)=\left(-\left(q_{1}^{(2)}p_{1}^{(2)}+q_{2}^{(2)}p_{2}^{(2)}-\alpha_{3}+\eta\right)p_{1}^{(2)},q_{2}^{(2)}p_{1}^{(2)},\dfrac{1}{p_{1}^{(2)}},\dfrac{p_{2}^{(2)}}{p_{1}^{(2)}}\right),$
$\left(\widehat{x}_{3}^{[21]},\widehat{y}_{3}^{[21]},\widehat{z}_{3}^{[21]},\widehat{w}_{3}^{[21]}\right)=r_{0}\left(x_{3}^{[21]},y_{3}^{[21]},z_{3}^{[21]},w_{3}^{[21]}\right),$
$C_{3}^{(22)}:$ $\left(x_{3}^{[22]},y_{3}^{[22]},z_{3}^{[22]},w_{3}^{[22]}\right)=\left(q_{1}^{(2)}p_{2}^{(2)},-\left(q_{1}^{(2)}p_{1}^{(2)}+q_{2}^{(2)}p_{2}^{(2)}-\alpha_{3}+\eta\right)p_{2}^{(2)},\dfrac{p_{1}^{(2)}}{p_{2}^{(2)}},\dfrac{1}{p_{2}^{(2)}}\right),$
$\left(\widehat{x}_{3}^{[22]},\widehat{y}_{3}^{[22]},\widehat{z}_{3}^{[22]},\widehat{w}_{3}^{[22]}\right)=r_{0}\left(x_{3}^{[22]},y_{3}^{[22]},z_{3}^{[22]},w_{3}^{[22]}\right)$
for $C_{3}^{0}$. Namely, the system on each coordinate $\left(x_{j}^{[kl]},y_{j}^{[kl]},z_{j}^{[kl]},w_{j}^{[kl]}\right)$ or $\left(\widehat{x}_{j}^{[kl]},\widehat{y}_{j}^{[kl]},\widehat{z}_{j}^{[kl]},\widehat{w}_{j}^{[kl]}\right)$ is described as a polynomial Hamiltonian system.
Here we use the notation
$r_{2}\left(x_{3}^{[01]},y_{3}^{[01]},z_{3}^{[01]},w_{3}^{[01]}\right)$
$=\left(r_{2}\left(-(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)p_{1}\right),r_{2}(q_{2}p_{1}),r_{2}\left(\dfrac{1}{p_{1}}\right),r_{2}\left(\dfrac{p_{2}}{p_{1}}\right)\right)$
$=\bigg(-\frac{(q_{1}p_{1}+q_{2}p_{2}-\alpha_{5}+\eta)\{(q_{1}-q_{2})p_{1}-\alpha_{2}\}}{q_{1}-q_{2}},\frac{q_{2}\{(q_{1}-q_{2})p_{1}-\alpha_{2}\}}{q_{1}-q_{2}},$
$\quad\qquad\frac{q_{1}-q_{2}}{(q_{1}-q_{2})p_{1}-\alpha_{2}},\frac{(q_{1}-q_{2})p_{2}+\alpha_{2}}{(q_{1}-q_{2})p_{1}-\alpha_{2}}\bigg)$
for the sake of simplicity. This proposition is shown in two ways. One is the blowing up method and another is that in terms of formal power series. Note that the latter method will be used in the next section to resolve $P^{0}$. Because we can prove in the same way for all of the accessible singularities, we consider only
$C_{1}^{0}$ $=C_{1}^{(01)}\cup C_{3}^{(01)}\cup C_{3}^{(02)}\cup C_{2}^{(12)}$
$=\left\{q_{1}-q_{2}=p_{1}^{(01)}=p_{2}^{(01)}+1=0\right\}\cup\left\{q_{1}=q_{2}=p_{1}^{(01)}=0\right\}$
$\quad\cup\left\{q_{1}=q_{2}=p_{2}^{(02)}=0\right\}\cup\left\{q_{2}^{(1)}-1=p_{1}^{(12)}=p_{2}^{(12)}=0\right\}$
here.
5.1 Blowing up
Let $\left(u_{1i},u_{2i},v_{1i},v_{2i}\right)$, $\left(u^{\prime}_{1i},u^{\prime}_{2i},v^{\prime}_{1i},v^{\prime}_{2i}\right)$ and $\left(u^{\prime\prime}_{1i},u^{\prime\prime}_{2i},v^{\prime\prime}_{1i},v^{\prime\prime}_{2i}\right)\ (i=1,2,3,4)$ be affine coordinates of $\mathbb{C}^{4}$ defined by
$q_{1}-q_{2}=u_{11}v_{11},\quad q_{2}=u_{21},\quad p_{1}^{(01)}=v_{11},\quad p_{2}^{(01)}+1=v_{11}v_{21},$
$q_{1}-q_{2}=u_{11}^{\prime},\quad q_{2}=u_{21}^{\prime},\quad p_{1}^{(01)}=u_{11}^{\prime}v_{11}^{\prime},\quad p_{2}^{(01)}+1=u_{11}^{\prime}v_{21}^{\prime},$
$q_{1}-q_{2}=u_{11}^{\prime\prime}v_{21}^{\prime\prime},\quad q_{2}=u_{21}^{\prime\prime},\quad p_{1}^{(01)}=v_{11}^{\prime\prime}v_{21}^{\prime\prime},\quad p_{2}^{(01)}+1=v_{21}^{\prime\prime}$
for $C_{1}^{(01)}$,
$\begin{split}&q_{1}=u_{12}v_{12},\quad q_{2}=u_{22}v_{12},\quad p_{1}^{(01)}=v_{12},\quad p_{2}^{(01)}=v_{22},\\ &q_{1}=u_{12}^{\prime},\quad q_{2}=u_{12}^{\prime}u_{22}^{\prime},\quad p_{1}^{(01)}=u_{12}^{\prime}v_{12}^{\prime},\quad p_{2}^{(01)}=v_{22}^{\prime},\\ &q_{1}=u_{12}^{\prime\prime}u_{22}^{\prime\prime},\quad q_{2}=u_{22}^{\prime\prime},\quad p_{1}^{(01)}=u_{22}^{\prime\prime}v_{12}^{\prime\prime},\quad p_{2}^{(01)}=v_{22}^{\prime\prime}\end{split}$ (5.1)
for $C_{3}^{(01)}$,
$q_{1}=u_{13}v_{23},\quad q_{2}=u_{23}v_{23},\quad p_{1}^{(02)}=v_{13},\quad p_{2}^{(02)}=v_{23},$
$q_{1}=u_{13}^{\prime},\quad q_{2}=u_{13}^{\prime}u_{23}^{\prime},\quad p_{1}^{(02)}=v_{13}^{\prime},\quad p_{2}^{(02)}=u_{13}^{\prime}v_{23}^{\prime},$
$q_{1}=u_{13}^{\prime\prime}u_{23}^{\prime\prime},\quad q_{2}=u_{23}^{\prime\prime},\quad p_{1}^{(02)}=v_{13}^{\prime\prime},\quad p_{2}^{(02)}=u_{23}^{\prime\prime}v_{23}^{\prime\prime}$
for $C_{3}^{(02)}$ and
$q_{1}^{(1)}=u_{14},\quad q_{2}^{(1)}-1=u_{24}v_{24},\quad p_{1}^{(12)}=v_{14}v_{24},\quad p_{2}^{(12)}=v_{24},$
$q_{1}^{(1)}=u_{14}^{\prime},\quad q_{2}^{(1)}-1=u_{24}^{\prime},\quad p_{1}^{(12)}=u_{24}^{\prime}v_{14}^{\prime},\quad p_{2}^{(12)}=u_{24}^{\prime}v_{24}^{\prime},$
$q_{1}^{(1)}=u_{14}^{\prime\prime},\quad q_{2}^{(1)}-1=u_{24}^{\prime\prime}v_{14}^{\prime\prime},\quad p_{1}^{(12)}=v_{14}^{\prime\prime},\quad p_{2}^{(12)}=v_{14}^{\prime\prime}v_{24}^{\prime\prime}$
for $C_{2}^{(12)}$. On these coordinates, the accessible singularity $C_{1}^{0}$ corresponds to
$\begin{split}D_{1}^{1}&=\left\{v_{11}=0\right\}\cup\left\{u_{11}^{\prime}=0\right\}\cup\left\{v_{21}^{\prime\prime}=0\right\}\cup\left\{v_{12}=0\right\}\cup\left\{u_{12}^{\prime}=0\right\}\cup\left\{u_{22}^{\prime\prime}=0\right\}\\ &\quad\cup\left\{v_{23}=0\right\}\cup\left\{u_{13}^{\prime}=0\right\}\cup\left\{u_{23}^{\prime\prime}=0\right\}\cup\left\{v_{24}=0\right\}\cup\left\{u_{24}^{\prime}=0\right\}\cup\left\{v_{14}^{\prime\prime}=0\right\}.\end{split}$ (5.2)
The union $D_{1}^{1}$ is connected and called an exceptional divisor. We consider the accessible singularity $C_{3}^{(01)}$. System (3.6) on the coordinate $(u_{12},u_{22},v_{12},v_{22})$ is described as
$\begin{split}&t(t-1)v_{12}du_{12}-U_{12}dt=0,\\ &t(t-1)v_{12}du_{22}-U_{22}dt=0,\\ &t(t-1)dv_{12}-V_{12}dt=0,\\ &t(t-1)dv_{22}-V_{22}dt=0,\end{split}$ (5.3)
where $U_{12},U_{22},V_{12},V_{22}\in\mathbb{C}[u_{12},u_{22},v_{12},v_{22},t]$. The singularity and the accessible singularity of system (5.3) on $D_{1}^{1}$ are given by
$\left\{v_{12}=0\right\}$